1. What is the theoretical yield of silver chloride (AgCl) if you begin the reaction with 0.2593 g of NaCl?

AgNO3(aq)+NaCl(s)---->AgCl(s)+NaNO3(aq)

2. Based on your above calculated theoretical yield, what is the percent yield of AgCl if you produce 0.4829 g

experimentally?
Answer #1

AgNO3 + Nacl - Agel & No Noz mol Nach = 0. 2593 1.926810-3 mol 58.45 = 4.436810-3 mol Agcl = man = man = 4.436x10 -3 mol (4.4